We have a mixture that is 20 mol% propane, 35 mol% n-butane, and 45 mol% nhexane. If a flash drum operates at 400 kPa, what is the highest temperature at which the flash drum can operate and still have vapor and liquid present? Use the DePriester chart for equilibrium
Answer #1

For the heating step, we calculate mean heat capacities between 298.15298.15 K 298.15 K and 460 K bYEqs. (4.7) and (5.17) with data from Table 4.1. The results in J mor l K- I are summarized as follows: 9.405 N, CO2 CPmh CPml H2O(b) Standard reaction at 298.15 K 41.649 41.377 N2 34.153 34.106 29.381 29.360 02 30.473 30.4051 mol 2 mol 0.5 molCO2 0, H,O (c) Mix at 298.15 K We multiply each individual heat capacity by the number of moles of that species in the flue gas and sum over all species. This gives total mean heat capacities for the 12.905 mol of mixture: CL = 401.520 and C~nu = 400.922 J K-I Then I1Hd = Ckh(T2 - Td = (401.520)(460 - 298.15) 12.905 mol = 64,986J ftue gas 460K I1Sd = C~ In T2 = 400.922 In 460 = 173.852 J K- I TI 298.15Figure 16.2 Calculation path for combustion process of Example 16.5. m. Step a: For unmixing the entering air